# Hai-Long Sun1,2,3   Da-Wei Zhou1,2  Yang Li3  Shiyin Lu3  Chao Yi1,2 Qing-Guo Chen3  Zhao Xu3  Weihua Luo3  Kaifu Zhang3  De-Chuan Zhan1,2  Han-Jia Ye1,2

### v2

@@ Appendix B Training Details
Table 5: Training hyperparameters.
@@
Appendix C Extended Experiments
In this section, we further provide more experiments and ablation studies to validate the generality and capability of Parrot across various tasks. Additionally, we present more training details about Figure 1 to offer a clearer understanding for readers.
C.1 Bilingual Evaluation on LLaVA-Bench
VisCPM [24] extends the LLaVA-Bench dataset to the Chinese version for bilingual evaluation. To comprehensively compare Parrot with other multilingual models, we conduct experiments on this benchmark. Due to the deprecation of the GPT-4-0314 version by OpenAI, we test Parrot in LLaVA-Bench following the version of GPT-4-1106-preview for comparison. As shown in Table 6, Parrot not only demonstrates exceptional ability in the English version of this benchmark but also presents competitive performance in the Chinese version.
Notably, as shown in Table 7, VisCPM uses 140M English data and 1M Chinese data to train the model, while Qwen-VL-Chat uses 1.1B English data and 300M Chinese data, whereas Parrot only utilizes approximately 2M data in total. Despite using less than 1% of the training data, Parrot achieves remarkable performance in both the English and Chinese versions on LLaVA-Bench. Owing to the architecture we proposed, significant improvement in the model’s multilingual capability can be achieved with minimal data usage.
Figure 8: An example of circular evaluation strategy.
Table 6: Experimental results on LLaVA Test Set accessed by GPT-4. Con: Conversation, DD: Detailed Description, CR: Complex Reasoning, AVG: the average score of three tasks. The best/second best results are marked in bold and underlined, respectively. The symbol denotes that the data are judged following the version of GPT-4-1106-preview because the GPT-4-0314 version is deprecated by OpenAI.
Model LLM Backbone English Chinese
Con DD CR AVG Con DD CR AVG
English Model MiniGPT-4 Vicuna-13B 65.0 67.3 76.6 69.7 - - - -
InstructBLIP Vicuna-13B 81.9 68.0 91.2 80.5 - - - -
LLaVA Vicuna-13B 89.5 70.4 96.2 85.6 - - - -
VisualGLM ChatGLM-6B 62.4 63.0 80.6 68.7 76.6 87.8 83.6 82.7
Qwen-VL-Chat Qwen-7B 82.4 76.9 91.9 83.8 82.3 93.4 89.5 88.2
VisCPM-Balance CPM-Bee-10B 75.5 64.7 91.3 77.3 85.4 81.4 96.6 88.0
Multilingual
Model
Parrot Qwen1.5-7B 82.5 71.0 89.3 81.1 82.1 88.6 92.3 87.7
Table 7: Comparison of vision encoders, LLMs, and training data.
Model vision encoder LLM Training Data
mPLUG-Owl ViT-L/14 (0.3B) BLOOMZ-7B -
VisualGLM Q-Former (1.6B) ChatGLM-6B English: 300M; Chinese 30M
Qwen-VL-Chat ViT-bigG (1.9B) Qwen-7B English: 1.1B; Chinese: 300M
VisCPM Muffin (0.7B) CPM-Bee-10B English: 140M; Chinese: 1M
Parrot ViT-L/14 (0.3B) Qwen1.5-Chat-7B English: 1.8M; Chinese: 71K
C.2 Radar Charts on MMBench and MMMB
For a more intuitive demonstration of the multilingual capabilities of Parrot, we present radar charts for the multilingual MMBench and MMMB benchmarks. As depicted in Figure LABEL:figure:radar_mmbench and Figure LABEL:figure:radar_mmmb, our proposed method Parrotexhibits significantly better performance compared to other models.
C.3 More Experimental Details about Different Backbones
In this section, we provide detailed information to explain Figure 1. Firstly, to ensure a fair comparison between the OpenAI-CLIP-based model and the Chinese-CLIP-based model, we train distinct models using the same training data as LLaVA, as shown in Table 4(a). The hyperparameters are listed in Table 5 without the MoE hyperparameters. As depicted in Figure 1, the OpenAI-CLIP-based model struggles to generate Chinese outputs when given Chinese prompts due to the English-centric training data. In contrast, despite the extremely scarce amount of Chinese training data, the Chinese-CLIP-based model naturally acquires zero-shot capability to understand, process, and generate Chinese texts. Furthermore, we compare both models on MMBench-CN and MMMB-zh to evaluate their Chinese capability. As shown in Table 8, the performance of the Chinese-CLIP-based model is significantly higher than that of the OpenAI-CLIP-based model. On the other hand, we empirically find that different LLMs have a significant impact on performance. Qwen [5] demonstrates superior Chinese capability compared to Vicuna [15], yet its English capability remains competitive.
Method Vision encoder LLM MMB-EN MMB-CN MMMB-en MMMB-zh
LLaVA OpenAI-CLIP ViT-L/14 Vicuna 7B 65.4 58.3 67.1 58.8
LLaVA OpenAI-CLIP ViT-L/14 Qwen1.5-Chat 7B 68.8 66.4 68.2 62.4
LLaVA Chinese-CLIP ViT-L/14 Qwen1.5-Chat 7B 68.1 68.3 67.6 66.1
Parrot OpenAI-CLIP ViT-L/14 Qwen1.5-Chat 7B 70.7 70.4 70.0 68.1
Appendix D Broader Impact and Limitations
Broader Impact. Parrot leveraging MoE to enhance multilingual alignment presents a positive social impact by promoting linguistic diversity and inclusivity. To address the challenge of the imbalanced language data in SFT datasets and improve non-English visual tokens alignment, this approach contributes to breaking language barriers and facilitating cross-cultural communication, thereby fostering understanding and collaboration across diverse linguistic communities. Additionally, the creation of the Massive Multilingual Multimodal Benchmark (MMMB) fills a crucial gap in evaluating multilingual capabilities, enabling researchers to assess and improve upon models’ performance across different languages and cultures. However, it’s crucial to acknowledge potential negative social impacts, such as the risk of hallucination. This could potentially result in the model generating misleading or incorrect information, which is a common challenge observed in MLLMs.
Limitations. Despite advancements, MLLMs may still exhibit limitations in accurately understanding and responding to complex language-specific contexts, leading to misinformation or misinterpretation of multilingual inputs. On the other hand, due to the visual component of Parrot being based on CLIP, there are inherent limitations in its ability to process high-resolution images, resulting in the inability to recognize extremely detailed content in some images. Hence, enhancing Parrot’s ability to handle high-resolution processing will be part of future work.
Appendix E More Visualization Results
@@
In this section, we include additional visualization results between users’ questions and Parrot’s responses using multiple languages. These pictures are selected from LLaVA [38] and CuMo [30]. As depicted in Figures LABEL:figure:case_en, LABEL:figure:case_zh, LABEL:figure:case_pt, LABEL:figure:case_ar, LABEL:figure:case_tr and LABEL:figure:case_ru, it is evident that Parrotpossesses superior multilingual capabilities for understanding, processing, and generating multilingual texts. In certain specific cases, Parrot may also experience hallucinations. As depicted in the upper case of Figure LABEL:figure:case_en, it misidentifies Xiaomi SU7 as a Porsche Taycan.

### v3

@@ D.3 Implementation Details of Parrot
Table 13: The detailed training hyperparameters.
@@
D.4 Analysis of the Translation-based Baseline
There is a naive baseline where we first translate the question into English and then translate the English answer back to the target language. On the one hand, our experimental setting follows recent work in multilingual and multimodal large language models (Hu et al., 2023; Zhang et al., 2024a; Hinck et al., 2024), where such a naive baseline has not been commonly considered. While the translation-based approach could be a straightforward alternative, it faces some significant challenges.
First, it is highly susceptible to translation noise, particularly issues related to polysemy and meaning ambiguity between languages. Moreover, our benchmark includes a substantial number of cultural-specific questions, which require deep cultural context knowledge that translation alone cannot effectively capture. In practical use, adding an additional translation step would also introduce extra overhead, increasing both the time and computational cost.
Despite these challenges, we acknowledge the importance of evaluating this baseline and conducting experiments to assess the performance of this translation-based baseline by using the Google Translation API. As shown in the Table 10, the results reveal a "seesaw effect"—while the naive baseline shows some improvements in certain languages, such as Chinese, it leads to performance degradation in others, such as Russian and Portuguese. This highlights the difficulty of addressing multilingualism and multimodal tasks solely through translation.
Appendix E Discussion
Parrot is a novel approach that leverages textual guidance to align visual tokens at the language level, enabling the conversion of English-biased visual embeddings into language-specific ones through an MoE module. In future work, we plan to incorporate more culture-related samples in various languages. This will enhance the representation of diverse cultural contexts and ensure that our benchmark accurately reflects the complexities of multilingual interactions. Additionally, we will focus on developing tasks that not only assess linguistic capabilities but also evaluate cultural nuances, which are crucial for effective communication in multilingual settings. By doing so, we aim to provide a more comprehensive evaluation of multilingual models and their performance across different cultural backgrounds.
Figure 11: Several culture-related samples in different languages.
Appendix F More Visualization Results
@@
In this section, we include additional visualization results between users’ questions and Parrot’s responses using multiple languages. These pictures are selected from LLaVA (Liu et al., 2023b) and CuMo (Li et al., 2024c). As depicted in Figures LABEL:figure:case_en, LABEL:figure:case_zh, LABEL:figure:case_pt, LABEL:figure:case_ar, LABEL:figure:case_tr and LABEL:figure:case_ru, it is evident that Parrotpossesses superior multilingual capabilities for understanding, processing, and generating multilingual texts. In certain specific cases, Parrot may also experience hallucinations. As depicted in the upper case of Figure LABEL:figure:case_en, it misidentifies Xiaomi SU7 as a Porsche Taycan.
Table 14: Ablation study on monolingual fine-tuning dataset in MMMB benchmark. The table shows an effect of performance on six languages when using fine-tuning data from different languages. Models with 7B parameters are used for this ablation.
Dataset English Chinese Portuguese Arabic Turkish Russian
LLaVA-1.5-finetune 72.69 67.60 65.61 57.72 48.30 63.80
+ zh 71k 69.18 69.06 63.92 58.13 48.95 63.63
+ pt 14k 69.94 68.83 65.67 58.65 51.11 63.04
+ ar 12k 70.47 68.36 64.39 60.79 51.11 63.16
+ tr 17k 70.82 69.01 64.85 60.76 60.70 64.39
+ ru 14k 69.59 68.07 64.27 60.35 53.92 64.15
+ zh pt ar tr ru 70.00 68.13 67.31 62.69 58.01 66.26